Transform each graph as specified below. (a) The graph of y=f(x) is shown. Graph y=f(x). (b) The graph of y=g(x) is shown. GrPlease graph well
Answer #1

@ Graph of yetens is shown, we need to fine the graph of y=f(tm). 1 2 3 4 DBE - 4 2 -0 We can clearly observe that OA and ABNow, our functions becomes y = f(x) = -24 , 2450,2) E3X-10, 2€ [2, 4]. Now Y= 44x) = 2x42:-1, I NE D.2 » 3: -x , x < [ 4) ny:y Jews is given, Po graph y: 2 gen. In this we clearly observe that the valles of a remains const the same ile the domein rem